Grade 12th passPhysical Chemistry

Why specific conductance decreases with increase in dilution?

Ankit Dash
7 Years agoGrade 12th pass

1 Answer

Arun
7 Years ago
Dear Ankit
 

Specific conductivity is the conductance by an electrolytic solution of unit volume kept between 2 platinum electrodes of unit length .Thus it  depends upon concentration of  the solution.

As volume increases on dilution, therefore no. of ions per unit volume carrying the current decreases. So , specific conductivity too decreases.

While molar conductivity is the conductivity per molar solution. In other words , conductivity of 1mole of unit volume of electrolytic solution. Therefore   on dilution molar conductivity increases.

k = l/AR

λm = k/M = kV/n